OpenStudy (anonymous):

Express in lowest terms.. x+x^4/x^2-x+1

OpenStudy (anonymous):

(x+x^4)/(x^2-x+1)=x(1+x^3 )/(x^2-x+1)=x(1+x)(1-x+x^2 )/(1-x+x^2 )=x(1+x)